Bugs in Computers & Typesetting A, 1996 1

This is a list of all corrections made to Computers & Typesetting, Volume A (also known as The TEXbook), between 1992 and the publication of the final printed version of that book in September 1996.

Page A31, line 8 (3/6/95)

TEX begins its error messages with ‘!’, and it shows what it was reading at the

Page A46, line 8 (1/22/95)

out for the occasional times when the adjacent characters aa, ae, and o/ should not be

Page A282, lines 11 and 12 (4/18/96)

unadorned ⟨box⟩ command, except that the new box being appended to the vertical list is also shifted left or right by the specified amount.

Page A285, lines 15–17 from the bottom (4/18/96)

\raise⟨dimen⟩⟨box⟩, \lower⟨dimen⟩⟨box⟩. This acts just like an unadorned ⟨box⟩ command, except that the new box being appended to the horizontal list is also shifted up or down by the specified amount.

Page A290, lines 4–6 from the bottom (4/18/96)

\raise⟨dimen⟩⟨box⟩, \lower⟨dimen⟩⟨box⟩. This acts just like an unadorned ⟨box⟩ command, except that the new box being put into the nucleus is also shifted up or down by the specified amount.

Page A331, bottom two lines (6/25/93)

if you know that the enclosing box is sufficiently small; and \leaders\vrule\vfill works fine in vertical mode.

Page A354, lines 19–22 (3/5/95)

\def\sett@b{\ifx\next\+\def\nxt{\afterassignment\s@tt@b\let\nxt}%
  \else\let\nxt=\s@tcols\fi
  \let\next=\relax \nxt} % turn off \outerness
\def\s@tt@b{\let\nxt=\relax \us@false\m@ketabbox}

Page A356, lines 13–20 from the bottom (3/5/95)

\def\oalign#1{\leavevmode\vtop{\baselineskip0pt \lineskip.25ex
  \ialign{##\crcr#1\crcr}}} \def\o@lign{\lineskiplimit=0pt \oalign}
\def\ooalign{\lineskiplimit=-\maxdimen \oalign} % chars over each other
\def\sh@ft#1{\dimen0=.00#1ex \multiply\dimen0 by\fontdimen1\font
  \kern-.0156\dimen0} % compensate for slant in lowered accents
\def\d#1{{\o@lign{\relax#1\crcr\hidewidth\sh@ft{10}.\hidewidth}}}
\def\b#1{{\o@lign{\relax#1\crcr\hidewidth\sh@ft{29}%
    \vbox to.2ex{\hbox{\char'26}\vss}\hidewidth}}}

Page A357, lines 7–12 (8/1/95)

\def\rightarrowfill{$\m@th \smash- \mkern-7mu
\cleaders\hbox{$\mkern-2mu \smash- \mkern-2mu$}\hfill
\mkern-7mu \mathord\rightarrow$}
\def\leftarrowfill{$\m@th \mathord\leftarrow \mkern-7mu
\cleaders\hbox{$\mkern-2mu \smash- \mkern-2mu$}\hfill
\mkern-7mu \smash-$}

Page A357, lines 16–20 (6/25/93)

\setbox0=\hbox{$\braceld$}%
\bracelu\leaders\vrule height\ht0 depth0pt\hfill\bracerd
\braceld\leaders\vrule height\ht0 depth0pt\hfill\braceru$}
\def\downbracefill{$\m@th
\setbox0=\hbox{$\braceld$}%
\braceld\leaders\vrule height\ht0 depth0pt\hfill\braceru
\bracelu\leaders\vrule height\ht0 depth0pt\hfill\bracerd$}

[Also delete lines 21 and 22, as the usage is no longer restricted.]

Page A359, line 25 (3/5/95)

\def\skew#1#2#3{{\muskip0=#1mu \mkern.5\muskip0
  #2{\mkern-.5\muskip0{#3}\mkern.5\muskip0}\mkern-.5\muskip0}{}}

Page A360, line 5 from the bottom (3/5/95)

\def\@vereq#1#2{\lower.5pt\vbox{\lineskiplimit\maxdimen \lineskip-.5pt

Page A361, lines 19 and 20 (3/5/95)

\def\bmod{\nonscript\mskip-\medmuskip \mkern5mu
\mathbin{\rm mod} \penalty900 \mkern5mu \nonscript\mskip-\medmuskip}

Page A362, lines 14–18 (3/5/95)

\everycr{\noalign{\ifdt@p \global\dt@pfalse \ifdim\prevdepth>-1000pt
    \vskip-\lineskiplimit \vskip\normallineskiplimit \fi
    \else \penalty\interdisplaylinepenalty \fi}}}
\def\@lign{\tabskip=0pt\everycr={}} % restore inside \displ@y
\def\displaylines#1{\displ@y \tabskip=0pt

Page A363, lines 8–9 from the bottom (12/8/89)

\if@mid \dimen@=\ht0 \advance\dimen@ by\dp\z@ \advance\dimen@ by12\p@
  \advance\dimen@ by\pagetotal \advance\dimen@ by-\pageshrink

Page A364, line 5 from the bottom (4/11/96)

\def\fmtname{plain}
\def\fmtversion{3.141592} % identifies the current format

Page A374, line 3 (3/7/95)

        \begingroup\aftergroup\def\aftergroup\asts\aftergroup{

Page A410, line 23 (4/18/96)

  Tema con variazione \(su un tema differente)

(This implies a corresponding correction to page 411.)

Page A451, line 16 (8/8/93)

But when plain TEX is tried on the name of a famous Welsh village,

Page A462, right column (3/5/95)

\cong ( ≅ ), 151, \underline{360}, 436.

Page A463, right column (6/25/93)

direct sum, see \oplus.

Page A464, left column (6/25/93)

\downbracefill ( ), 225–226, \underline{357}.

Page A483, lines 15–21 (4/29/97)

       email: {\tt TUG@tug.org}
       internet: {\tt http://www.tug.org/}
       }

Don’t delay, subscribe today! That address again is TEX Users Group

       email: TUG@tug.org
       internet: http://www.tug.org/